What is the density of atmosphere, assuming N2 : 79% and O2 : 21% ?
My chemistry instructor gave us this question to answer, and I feel as if its a trick question.
Might it have something to do with the ideal gas law?
1
Expert's answer
2019-04-15T07:54:30-0400

In 100g of air:

"n(O2)=\dfrac{21g}{32g/mol}=0.66 moles"

"n(N2)=\dfrac{79g}{28g/mol}=2.82moles"

"0.66moles+2.82moles=3.48 moles"

"V=\dfrac{nRT}{P}=\dfrac{3.48*0.0821*273K}{1atm}=78L"

"density=\dfrac{m}{V}=\dfrac{100g}{78L}=1.282g/L=1.282kg/m^{3}"